OpenStudy (anonymous):

Find the product: (2-6i)(7-5i)

OpenStudy (anonymous):

I know the answer is 16-52i,but have no idea why

OpenStudy (anonymous):

I originally got 14-52i

OpenStudy (anonymous):

*the answer is -16-52i

OpenStudy (anonymous):

But I want to know why

OpenStudy (jdoe0001):

\(\bf (2-6i)(7-5i) \implies 14-10i-42i+30i^2\\ \quad \\ \quad \\ 14+30i^2-52i\\ \textit{keep in mind that } i^2 = \sqrt{ (-1)}\times \sqrt{ (-1)} \implies \sqrt{ (-1)^2} \implies -1\)

OpenStudy (anonymous):

I still don't understand why that would equal -16-52i

OpenStudy (jdoe0001):

\(\bf (2-6i)(7-5i) \implies 14-10i-42i+30i^2\\ \quad \\ \quad \\ 14+30i^2-52i \\ \textit{keep in mind that } i^2 = \sqrt{ (-1)}\times \sqrt{ (-1)} \implies \sqrt{ (-1)^2} \implies -1\\\quad \\ 14+30i^2-52i \implies 14+30(-1)-52i \)

OpenStudy (jdoe0001):

see it now?

OpenStudy (jdoe0001):

\(\bf 14+30i^2-52i \implies 14+30(-1)-52i \implies 14-30-52i\)

OpenStudy (anonymous):

yes. Thank you. Not to bother you, but if we have (-6+i)(-6-i), why would that equal 37 rather than 37+6i?

OpenStudy (jdoe0001):

\(\bf (-6+i)(-6-i) \implies 36\cancel{+6i}\cancel{-6i}-i^2 \implies 36-i^2\\ \textit{keep in mind that } i^2 = \sqrt{ (-1)}\times \sqrt{ (-1)} \implies \sqrt{ (-1)^2} \implies -1\\ \quad \\ \quad \\ 36-i^2 \implies 36-(-1) \implies 37\)

OpenStudy (anonymous):

But I FOILd it and got 36+6i+1 and that's why I have 37+6i rather than just 37.

OpenStudy (jdoe0001):

\(\bf (-6+i)(-6-i) \\ \quad \\ \quad \\ -6 \times -6 = 6\\ -6 \times -i = 6i\\ \quad \\ \quad \\ +i \times -6 = -6i\\ +i \times -i = -i^2\)

OpenStudy (jdoe0001):

woops.... darn... got a typo...

OpenStudy (jdoe0001):

\(\bf (-6+i)(-6-i) \\ \quad \\ \quad \\ -6 \times -6 = 6^2\\ -6 \times -i = 6i\\ \quad \\ \quad \\ +i \times -6 = -6i\\ +i \times -i = -i^2\)

OpenStudy (jdoe0001):

in FOIL it should give you the same values anyhow